OpenStudy (anonymous):

"Calculus Lec1" Find the area of triangle enclosed by the axes and tangent to y=1/xPlease explain how to get the y-intercept algebraically not by symmetry. x=2x0 part.

OpenStudy (anonymous):

we have the equation y-y0 = {-1/(xo)^2} * x-xo put x=0 for y intercept to get y - yo = {-1/(xo)^2}* 0-xo -> y-yo = -1/(xo)^2 * -xo -> y - yo = 1/xo -> y - yo = yo as [yo = 1/xo] -> y = 2yo